Desired consumption, desired investment, and government spending in a closed economy are

Cd = 260 - 100r + 0.2Y

Id = 100 - 300r

G = 220

What value of the real interest rate clears the goods market when Y = 600? (Please show your work)
Answer #1

Answer

the equilibrium is at

Y=C+I+G

Y= 260 - 100r + 0.2Y+100 - 300r+220

0.8Y=580-400r

400r=580-0.8Y

Y=600

400r=580-0.8*600

r=100/400

r=0.25=25%

the interest rate is 25%